# bayesNMF: Fast Bayesian Poisson NMF with Automatically Learned Rank Applied to Mutational Signatures

### v2

@@ Abstract
Bayesian Non-Negative Matrix Factorization (NMF) is a method of interest in fields including genomics, neuroscience, and audio and image processing. Bayesian Poisson NMF is of particular importance for counts data, for example in cancer mutational signature analysis. However, MCMC methods for Bayesian Poisson NMF require a computationally intensive augmentation. Further, identifying latent rank is necessary, but commonly used heuristic approaches are slow and potentially subjective, while methods that learn rank automatically are unable to provide posterior uncertainties. In this paper, we introduce bayesNMF, a computationally efficient Gibbs sampler for Bayesian Poisson NMF. Metropolis-Hastings steps are used to avoid augmentation, where full conditionals from a Normal-likelihood NMF is used as geometry-informed, high-overlap proposals. We additionally define sparse Bayesian factor inclusion (SBFI) as a method to identify rank automatically while providing posterior uncertainty quantification. We provide an open-source R software package with all of the models and plotting capabilities demonstrated in this paper on GitHub at jennalandy/bayesNMF, and supplemental materials are available online. Although our applications focus on cancer mutational signatures, our software and results can be extended to any use of Bayesian Poisson NMF.
@@ 1 Introduction
Non-Negative Matrix Factorization (NMF) is a widely used unsupervised technique for dimension reduction and intuitive parts-based representation [Lee and Seung, 1999]. In cancer genomics, Poisson NMF can be applied to mutation counts to model mutational processes with latent factors, assuming that multiple processes contribute additively to tumor genomes [Alexandrov et al., 2013]. Bayesian Poisson NMF allows for uncertainty quantification and the integration of prior knowledge [Cemgil, 2009, Schmidt et al., 2009].
@@
The computational cost of MCMC methods (e.g., Gibbs samplers) for Bayesian NMF grows with the number of samples $G$ (columns), latent factors $N$, and variables $K$ (rows). With a Normal likelihood, conditional updates take the form of standard distributions, requiring $(K+G)\times N$ Gibbs updates per iteration, plus updates for residual variance [Schmidt et al., 2009]. In contrast, Bayesian NMF with a Poisson likelihood involves a data augmentation step, called Poisson augmentation, that decomposes each observed count into latent factor-specific counts, introducing $K\times G\times N$ auxiliary latent variables. This results in $(K+G+KG)\times N$ Gibbs updates [Cemgil, 2009], leading to a substantially higher computational cost. With growing sources of publicly available data for the analysis of mutational signatures [Consortium, 2010, Weinstein et al., 2013, Consortium, 2020] and a growing reference set of previously discovered signatures [Alexandrov et al., 2020], computational efficiency is a key consideration for work in this field.
@@
In addition, identifying latent rank is a fundamental aspect of any factor model. Most Bayesian NMF methods take a heuristic approach: they fit a separate model for each rank and optimize regularized metrics such as the Bayesian information criterion (BIC) [Rosales et al., 2016, Gori and Baez-Ortega, 2018] or determine the “elbow” of metrics such as mean squared error (MSE) or cosine similarity [Gori and Baez-Ortega, 2018, Islam et al., 2022]. This is a computationally intensive process and requires post hoc, and potentially subjective, model selection. Existing methods that automatically learn rank as part of the Bayesian model with automatic relevance determination (ARD) do so by optimizing the posterior numerically, which means they are fast but result in point estimates without posterior uncertainties [Kasar et al., 2015]. Another recent method, compressive NMF [Zito and Miller, 2024, CompNMF], is able to learn rank automatically while maintaining posterior uncertainty, but it does not ease the computational burden of Poisson augmentation.
@@
Table 1 summarizes the relative strengths of existing methods for Poisson Bayesian NMF and identifies the gap that this work addresses: we introduce a Gibbs sampler for Bayesian Poisson NMF which 1) is computationally efficient by avoiding Poisson augmentation, 2) automatically determines latent rank as a part of the Bayesian model, and 3) reports posterior uncertainties. We avoid the computationally intensive Poisson augmentation through Metropolis Hastings sampling with high-overlap, geometry-informed proposals derived from a Normal-likelihood Bayesian NMF. We show the accuracy and efficiency of this approach through simulation studies and apply it to 32 cancer types from the Pan-Cancer Analysis of Whole Genomes (PCAWG) database [Consortium, 2020]. We also provide an open-source R software package, bayesNMF, which is available on GitHub at jennalandy/bayesNMF.
@@ 2 Background
Table 1: Software for Bayesian NMF applied to de-novo mutational signature analysis: SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], SigneR [Rosales et al., 2016], SigFit [Gori and Baez-Ortega, 2018], and CompNMF [Zito and Miller, 2024]. An ideal software would learn rank automatically, report posterior uncertainty, and avoid computationally intensive Poisson augmentation.
@@ 2.1 Non-Negative Matrix Factorization (NMF) and Bayesian NMF
NMF is an unsupervised method to decompose non-negative data $M$ ($K$ variables by $G$ samples) into two lower-rank non-negative matrices, $P$ and $E$ [Paatero and Tapper, 1994, Lee and Seung, 1999]. NMF represents each observation of the data matrix, or column of $M$, as a linear combination of latent factors, or columns of $P$, with weights in $E$:
$M_{kg}$ $=\sum_{n=1}^{N}P_{kn}E_{ng},\hskip 11.38109pt\mbox{i.e}\;\;M=PE,\hskip 11.38109ptM\in\mathbb{R}_{\geq 0}^{K\times G},P\in\mathbb{R}_{\geq 0}^{K\times N},E\in\mathbb{R}_{\geq 0}^{N\times G}.$
NMF is typically fit by minimizing a reconstruction error, often the generalized Kullback-Leibler (KL) divergence between $M$ and $PE$ (Equation 1) or the Frobenius norm of MPEM-PE (Equation 2). For each, Lee and Seung [2000] adapt gradient descent by deriving a step size that preserves nonnegativity, resulting in multiplicative update algorithms.
$KL(M||PE)$ $=\sum_{kg}\left(M_{kg}log\frac{M_{kg}}{(PE)_{kg}}-M_{kg}+(PE)_{kg}\right)$ (1)
@@
Minimizing KL-divergence is equivalent to maximizing a Poisson likelihood on $M$ centered at $PE$ [Cemgil, 2009]. Similarly, minimizing the Frobenius norm is equivalent to maximizing a Normal likelihood on $M$ centered at $PE$ [Schmidt et al., 2009].
Bayesian approaches add priors on $P$ and $E$ [Cemgil, 2009, Schmidt et al., 2009]. To sample Bayesian Poisson NMF posterior distributions with standard MCMC methods, Poisson augmentation is necessary to achieve easily-sampled full conditional distributions of $P$ and $E$ [Cemgil, 2009]. This introduces latent variables $Z_{kng}$, defined as the part of the $M_{kg}$ count attributed to latent factor $n$. Each $Z_{kng}$ follows a Poisson distribution centered at $P_{kn}E_{ng}$ and satisfies $M_{kg}=\sum_{n}Z_{kng}$. This imposes $K\cdot N\cdot G$ additional Gibbs updates at each iteration, increasing the computational requirements of Bayesian Poisson NMF far beyond that of Bayesian Normal NMF.
@@ 2.2 Cancer Mutational Signatures
Cancer is driven by genetic changes, including somatic DNA mutations, which can be caused by multiple processes. Understanding which mutational process activities are reflected in a tumor genome can improve our understanding of cancer development and answer questions about subtyping, prognosis, and treatment [Nik-Zainal et al., 2012a, b, Alexandrov et al., 2013, Rosales et al., 2016]. A computationally-derived mutational signature models the specific patterns produced by a mutational process. Examples include DNA damage repair deficiencies [Zámborszky et al., 2017], deamination of 5-methylcytosine [Nik-Zainal et al., 2012a], tobacco smoke [Alexandrov et al., 2013, Nik-Zainal et al., 2015], and ultraviolet radiation [Nik-Zainal et al., 2015, Saini et al., 2016, Hayward et al., 2017]. The Catalog of Somatic Mutations in Cancer (COSMIC) [Tate et al., 2019] provides a high-confidence set of reference signatures. These are an important reference, but cannot be treated as ground truth, as they were estimated from data.
@@
Mutational signature analysis is a rapidly growing field driven by an increasing number of publicly available databases, including COSMIC [Tate et al., 2019], mSignatureDB [Huang et al., 2018], the Pan-Cancer Analysis of Whole Genomes (PCAWG) [Consortium, 2020], and the compendium of Mutational Signatures of Environmental Agents [Kucab et al., 2019]. With increased data sizes and availability, computational efficiency must be carefully considered when making modeling and implementation choices.
@@ 3.1 MH-within-Gibbs for Computational Efficiency
We introduce two efficient models for Poisson Bayesian NMF. The first uses a Truncated Normal (TN) prior on each element of $P$ and $E$:
@@
Priors: $\quad P_{kn}\sim\text{TN}(\mu^{P}_{kn},\sigma^{P}_{kn},0,\infty),\quad E_{ng}\sim\text{TN}(\mu^{E}_{ng},\sigma^{E}_{ng},0,\infty).$ (4)
@@
$f(P_{kn}^{*}|P_{-kn},E,M)$ $\propto p_{\text{TN}}(P_{kn}^{*})\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$ (5)
$\propto p_{\text{TN}}(P^{*}_{kn})e^{-\sum_{g}(P^{*}E)_{kg}}\prod_{g}(P^{*}E)_{kg}^{M_{kg}}.$
@@
$g(P^{*}_{kn}$ $|P_{-kn},E,M)\propto p_{\text{TN}}(P^{*}_{kn})\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$ (6)
$\sim\text{Truncated Normal}\left(P^{*}_{kn}|\mu=m(M,P,E,\theta),\sigma=s(M,P,E,\theta),0,\infty\right)$
@@
Instead of modeling variances $\sigma^{2}_{kg}$ as would be required in a Normal Bayesian NMF, we substitute it with $(PE)_{kg}$ in the proposal. This allows the proposal to reflect the Poisson mean-variance relationship of the target model without additional parameters. This only affects the proposal $g$, which needs to be a proper distribution but does not need to be an actual full conditional to maintain detailed balance.
@@
For a fixed rank, the Normal and Poisson Bayesian NMF models have identical MAP solutions (up to scaling and permutation symmetries of NMF [Lee and Seung, 1999, Donoho and Stodden, 2003]). Consequently, the Normal-based proposal captures the local curvature of the Poisson posterior near its mode, ensuring high overlap. This proposal is analogous to gradient-based methods (e.g., Metropolis-adjusted Langevin [Roberts and Tweedie, 1996] or Hamiltonian Monte Carlo [Neal et al., 2011]), but we do not compute gradients explicitly.
The second model we introduce uses an Exponential prior. The proposal still takes the form of a Truncated Normal distribution, though functions $m(\cdot)$ and $s(\cdot)$ change. In general, the Poisson+MH approach can only be used when the prior (which we assume is shared by the target and the proposal) leads to Normal Bayesian NMF full conditionals that are of known families, meaning Gamma priors cannot be used with this setup. The limitation is practical rather than fundamental—other proposal families could in principle accommodate Gamma priors but are not considered here.
@@ 3.2 Sparse Bayesian Factor Inclusion (SBFI) for Learning Rank
Previous Bayesian work by Grabski et al. [2023a, b] in the multi-study setting learns factor sharing patterns and total latent dimension using a binary factor inclusion matrix $A$. We adapt this approach to learn rank automatically in our single-study setting and introduce a new prior and hyperprior structure such that the induced prior on the expected rank is approximately uniform. We term this approach Bayesian factor inclusion (BFI).
@@
$\mathbb{E}[M_{kg}]=\sum_{n=1}^{N}P_{kn}A_{nn}E_{ng},\quad\quad\mbox{i.e.}\quad\quad\mathbb{E}[M]=PAE$
with latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$. When $A_{nn}=0$, hyperpriors of $P$ and $E$ retain some information on the excluded signature, reducing the extent of label switching.
Each factor inclusion indicator $A_{nn}$ has a Bernoulli prior. The prior probability of factor inclusion $q$ is determined by a random variable $R$, which represents the expected rank and has a uniform hyperprior between 0 and $N$. The value $q$ is bounded away from 0 and 1 to prevent the Gibbs sampler from getting stuck at the endpoint ranks 0 or $N$:
@@
In practice, we set $N$ to a large number. By learning the posterior distribution of $A$, we also learn the implied posterior distribution of the latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$.
Including $A$ in the model induces a highly multimodal posterior. We adapt the tempering approach used in Grabski et al. [2023b] for the Gibbs updates of $A$ and $R$, where the likelihood term is raised to a temperature power $\gamma$: $A_{nn}\sim p(A_{nn})p(M|A_{nn},...)^{\gamma}$ and $R\sim p(R)p(M|R,...)^{\gamma}$. The temperature $\gamma$ progresses from 0 to 1 over the course of a burn-in period. Early samples are from the prior alone, allowing better exploration of the parameter space, and once $\gamma=1$, samples are from the full conditional. No tempered samples are used for posterior inference.
We also introduce Sparse Bayesian factor inclusion (SBFI), which replaces the likelihood in the Gibbs update of $A_{nn}$ with $e^{-\frac{1}{2}\text{BIC}}$ to encourage sparsity in the latent rank:
BIC $=-2\cdot\text{log}\left[p_{\text{Poisson}}(M|A_{nn}=a,...)\right]+\text{log}(G)\cdot(K+G)N^{(a)}$
[…]
where $N^{(a)}=\sum_{n^{\prime}}A_{n^{\prime}n^{\prime}}$ when $A_{nn}=a$, $G$ is the sample size, and $(K+G)N^{(a)}$ is the number of functional parameters in $P$ and $E$. Expanding $e^{-\frac{1}{2}\text{BIC}}$ shows that, once $\gamma=1$, SBFI is equivalent to a Gibbs update under the following sparse prior on $A_{nn}$:
$p_{\text{SBFI}}(A_{nn}=1)$ $=\frac{qG^{-\frac{1}{2}(K+G)}}{qG^{-\frac{1}{2}(K+G)}+(1-q)}.$

### v3

@@ Abstract
Bayesian Poisson Non-Negative Matrix Factorization (NMF) is widely used to model count data, including in cancer mutational signature analysis. However, standard Gibbs samplers rely on computationally expensive Poisson augmentation, and current software implementations learn the latent rank either through slow and potentially subjective heuristic rank selection or with automatic approaches that do not report posterior uncertainty. In this paper, we introduce bayesNMF, an MH-within-Gibbs sampler to address both of these limitations. First, we define high-overlap proposals for Metropolis-Hastings sampling to remove the need for Poisson augmentation. Second, we define a BIC-based sparsity prior to learn rank automatically within the Bayesian formulation while allowing for posterior uncertainty quantification. We provide an open-source R software package with all of the models and plotting capabilities demonstrated in this paper on GitHub at jennalandy/bayesNMF. Although our applications focus on cancer mutational signatures, our software and results can be extended to any use of Bayesian Poisson NMF.
@@ 1 Introduction
Non-Negative Matrix Factorization (NMF) is a widely used unsupervised technique for dimension reduction and intuitive parts-based representation [Lee and Seung, 1999], and is typically formulated with either a Poisson or a Normal likelihood. In cancer genomics, Poisson NMF applied to mutation counts assumes that tumor genomes arise from multiple additive mutational processes, with the resulting latent factors termed mutational signatures [Alexandrov et al., 2013]. Within this framework, Bayesian NMF allows for integration of prior knowledge and posterior uncertainty quantification [Cemgil, 2009, Schmidt et al., 2009].
@@
The computational cost of a Gibbs sampler for Bayesian NMF grows with the number of samples $G$ (columns), latent factors $N$, and variables $K$ (rows). A Normal likelihood requires approximately $N(K+G)$ updates per iteration [Schmidt et al., 2009]. A Poisson likelihood, however, relies on Poisson augmentation for conditional updates to follow standard distributions. Each observed count is decomposed into latent factor-specific counts, introducing $NKG$ auxiliary parameters. This requires $N(K+G+KG)$ updates per iteration [Cemgil, 2009], resulting in a substantially higher computational cost. With growing publicly available data sources [Consortium, 2010, Weinstein et al., 2013, Consortium, 2020] and an expanding set of previously discovered signatures [Alexandrov et al., 2020], computational efficiency is a central constraint in large-scale mutational signature analysis.
@@
Further, identifying latent rank is a fundamental aspect of any factor model. Most Bayesian NMF methods take a heuristic approach, fitting a separate model for each rank and optimizing regularized metrics such as the Bayesian information criterion (BIC) [Rosales et al., 2016, Gori and Baez-Ortega, 2018] or determining the “elbow” of metrics such as MSE or cosine similarity [Gori and Baez-Ortega, 2018, Islam et al., 2022]. This process is computationally intensive and relies on post hoc and potentially subjective model selection.
Existing methods address subsets of these challenges, but trade off posterior uncertainty, computational efficiency, and rank learning in different ways (Table 1). SignatureAnalyzer [Kasar et al., 2015] uses automatic relevance determination (ARD) to learn rank as part of the Bayesian model. It avoids Poisson augmentation by optimizing the posterior numerically, but as a result, it provides only point estimates without posterior uncertainties. On the other hand, SigFit [Gori and Baez-Ortega, 2018] avoids Poisson augmentation by performing posterior inference in Stan [Carpenter et al., 2017], which implements Hamiltonian Monte Carlo [Neal and others, 2011], but selects rank heuristically across candidate models. Compressive NMF [Zito and Miller, 2024] learns rank automatically while maintaining posterior uncertainty, but relies on computationally expensive Poisson augmentation.
@@
To our knowledge, no method simultaneously 1) avoids Poisson augmentation, 2) learns rank within the Bayesian formulation, and 3) reports posterior uncertainty. This paper introduces two innovations to address this gap. First, we define high-overlap proposals for Metropolis-Hastings-within-Gibbs sampling to avoid Poisson augmentation. Second, we define a BIC-based sparsity prior to learn rank within the Bayesian model. The accuracy and efficiency of the combined approach are evaluated through simulation studies, and results are presented on 32 cancer types from the Pan-Cancer Analysis of Whole Genomes (PCAWG) database [Consortium, 2020]. Methods are implemented in an open-source R package, bayesNMF, available on GitHub at jennalandy/bayesNMF.
@@ 2 Background
Table 1: Software for Bayesian Poisson NMF applied to de-novo mutational signature analysis: SignatureAnalyzer [Kasar et al., 2015, Kim et al., 2016], SigneR [Rosales et al., 2016], SigFit [Gori and Baez-Ortega, 2018], and CompNMF [Zito and Miller, 2024]. An ideal software would automatically learn rank within the Bayesian formulation, report posterior uncertainty, and avoid computationally intensive Poisson augmentation.
@@ 2.1 Non-Negative Matrix Factorization (NMF) and Bayesian NMF
NMF is an unsupervised method to decompose non-negative data $M$ ($K$ variables as rows by $G$ samples as columns) into two lower-rank non-negative matrices, $P$ and $E$ [Paatero and Tapper, 1994, Lee and Seung, 1999]. NMF represents each observation of the data matrix (column of $M$) as a linear combination of latent factors (columns of $P$):
$M_{kg}$ $=\sum_{n=1}^{N}P_{kn}E_{ng},\hskip 11.38109pt\mbox{i.e.,}\;\;M=PE,\hskip 11.38109ptM\in\mathbb{R}_{\geq 0}^{K\times G},P\in\mathbb{R}_{\geq 0}^{K\times N},E\in\mathbb{R}_{\geq 0}^{N\times G}.$
NMF is typically fit by minimizing a reconstruction error, often the generalized Kullback-Leibler (KL) divergence (Equation 1) or the Frobenius norm (Equation 2).
$KL(M||PE)$ $=\sum_{kg}\left(M_{kg}\log\frac{M_{kg}}{(PE)_{kg}}-M_{kg}+(PE)_{kg}\right)$ (1)
@@
For each, Lee and Seung [2000] adapt gradient descent by deriving a step size that preserves nonnegativity, resulting in multiplicative update algorithms. These equate to probabilistic models: minimizing KL-divergence is equivalent to maximizing a Poisson likelihood on $M$ with mean $PE$ [Cemgil, 2009], and minimizing the Frobenius norm is equivalent to maximizing a Normal likelihood on $M$ with mean $PE$ [Schmidt et al., 2009].
Bayesian approaches place priors on $P$ and $E$ [Cemgil, 2009, Schmidt et al., 2009]. Sampling the posterior distribution of Bayesian Poisson NMF with a standard Gibbs sampler requires Poisson augmentation [Cemgil, 2009]. This introduces latent variables $Z_{kng}$, defined as the counts from $M_{kg}$ attributed to latent factor $n$. Each $Z_{kng}$ follows a Poisson distribution with mean $P_{kn}E_{ng}$ and satisfies $M_{kg}=\sum_{n}Z_{kng}$. These additional $KNG$ parameters increase the computational requirements of Bayesian Poisson NMF far beyond those of Bayesian Normal NMF.
@@ 2.2 Cancer Mutational Signatures
Cancer is driven by genetic changes, including somatic DNA mutations, which can be caused by multiple processes. Understanding which mutational process activities are reflected in a tumor genome can help answer questions about cancer development, subtyping, prognosis, and treatment [Nik-Zainal et al., 2012a, b, Alexandrov et al., 2013, Rosales et al., 2016]. A computationally-derived mutational signature models the specific patterns produced by a mutational process. Signatures have previously been linked to DNA damage repair deficiencies [Zámborszky et al., 2017], deamination of 5-methylcytosine [Nik-Zainal et al., 2012a], tobacco smoke [Alexandrov et al., 2013, Nik-Zainal et al., 2015], and ultraviolet radiation [Nik-Zainal et al., 2015, Saini et al., 2016, Hayward et al., 2017]. The Catalog of Somatic Mutations in Cancer (COSMIC) [Tate et al., 2019] provides a high-confidence set of reference signatures. These are an important reference, but cannot be treated as ground truth, as they were estimated from data.
@@
Mutational signature analysis is a rapidly growing field driven by an increasing number of publicly available databases, including COSMIC [Tate et al., 2019], mSignatureDB [Huang et al., 2018], PCAWG [Consortium, 2020], and the compendium of Mutational Signatures of Environmental Agents [Kucab et al., 2019]. With increased data sizes and availability, computational efficiency must be carefully considered when making modeling and implementation choices.
@@ 3.1 MH-within-Gibbs for Computational Efficiency
We introduce two efficient models for Poisson Bayesian NMF. The first uses a Truncated Normal prior on each element of $P$ and $E$:
@@
Priors: $\quad P_{kn}\sim\text{TruncNorm}(\mu^{P}_{kn},\sigma^{P}_{kn},0,\infty),\quad E_{ng}\sim\text{TruncNorm}(\mu^{E}_{ng},\sigma^{E}_{ng},0,\infty).$ (4)
@@
$f(P_{kn}^{*}|P_{-kn},E,M)$ $\propto p_{\text{TruncNorm}}(P_{kn}^{*})\prod_{g}p_{\text{Poisson}}\left(M_{kg}|\lambda=(P^{*}E)_{kg}\right)$ (5)
$\propto p_{\text{TruncNorm}}(P^{*}_{kn})e^{-\sum_{g}(P^{*}E)_{kg}}\prod_{g}(P^{*}E)_{kg}^{M_{kg}}.$
@@
$g(P^{*}_{kn}|P_{-kn},E,M)$ $\propto p_{\text{TruncNorm}}(P^{*}_{kn}|\theta)\prod_{g}p_{\text{Normal}}\left(M_{kg}|\mu=(P^{*}E)_{kg},\sigma^{2}=(PE)_{kg}\right)$ (6)
$=p_{\text{TruncNorm}}\left(P^{*}_{kn}|\mu=m(M,P,E,\theta),\sigma=s(M,P,E,\theta),0,\infty\right)$
@@
Instead of modeling variances $\sigma^{2}_{kg}$ as would be required in a Normal Bayesian NMF, we substitute $(PE)_{kg}$ into the proposal. This allows the proposal to reflect the Poisson mean-variance relationship of the target model without additional parameters. This affects only the proposal $g$, which needs to be a proper distribution in order to maintain detailed balance, but does not need to be an actual full conditional.
@@
For a fixed rank, the Normal and Poisson Bayesian NMF models have approximately the same MAP solutions (up to scaling and permutation symmetries of NMF [Lee and Seung, 1999, Donoho and Stodden, 2003]). Consequently, the Normal-based proposal captures the local curvature of the Poisson posterior near its mode, designed to have high overlap.
The second model we introduce places an Exponential prior on elements of $P$ and $E$. In order to adapt the Poisson+MH approach to this model, the proposal distribution is now based on a paired Normal likelihood NMF with Exponential priors. This proposal again takes the form of a Truncated Normal distribution but with redefined functions $m(\cdot)$ and $s(\cdot)$.
In general, the Poisson+MH approach can only be used when the prior (which we assume is shared by the target and the proposal) leads to Normal Bayesian NMF full conditionals that are of known families, meaning Gamma priors cannot be used with this setup. The limitation is practical rather than fundamental—other proposal families could, in principle, accommodate Gamma priors but are not considered here.
@@ 3.2 Sparse Bayesian Factor Inclusion (SBFI) for Learning Rank
Previous Bayesian work by Grabski et al. [2023, 2025] in the multi-study setting learns factor sharing patterns and total latent dimension using a binary factor inclusion matrix $A$. We adapt this approach to learn rank automatically in our single-study setting and introduce a new prior and hyperprior structure such that the induced prior on the expected rank is approximately uniform. We refer to this as Bayesian factor inclusion (BFI).
@@
$\mathbb{E}[M_{kg}]=\sum_{n=1}^{N}P_{kn}A_{nn}E_{ng},\quad\quad\mbox{i.e.,}\quad\quad\mathbb{E}[M]=PAE$
with a prespecified large $N$ and learned latent rank $N^{\prime}=\sum_{n=1}^{N}A_{nn}$. If $A$ is fixed as an identity matrix, this formulation is identical to that in Section 3.1 for latent rank $N$.
Each factor inclusion indicator $A_{nn}$ has a Bernoulli prior. The prior probability of factor inclusion $q$ is determined by a random variable $R$ representing the expected rank with a uniform hyperprior. The value $q$ is bounded away from 0 and 1 to prevent the Gibbs sampler from getting stuck at the endpoint ranks 0 or $N$:
@@
If $A_{nn}$ is sampled as 0, hyperpriors of $P$ and $E$ retain some information on the excluded signature, reducing the extent of label switching.
Including $A$ in the model induces a highly multimodal posterior. We adapt the tempering approach used in Grabski et al. [2025] for the Gibbs updates of $A$ and $R$, where the likelihood term is raised to a temperature power $\gamma$: $A_{nn}\sim p(A_{nn})p(M|A_{nn},...)^{\gamma}$ and $R\sim p(R)p(M|R,...)^{\gamma}$. The temperature $\gamma$ progresses from 0 to 1 over the course of a burn-in period. Early samples are from the prior alone, allowing better exploration of the parameter space, and once $\gamma=1$, samples are from the full conditional. No tempered samples are used for posterior inference.
We also introduce Sparse BFI (SBFI) with a sparse prior on $A_{nn}$:
$p$ $_{\text{SBFI}}(A_{nn}=1)=\frac{qG^{-\frac{1}{2}(K+G)}}{qG^{-\frac{1}{2}(K+G)}+(1-q)}$
$p$ $_{\text{SBFI}}(A_{nn}=a|...)\propto q^{a}(1-q)^{1-a}\cdot\left[p(M|A_{nn}=a,...)\cdot G^{-\frac{1}{2}(K+G)N^{(a)}}\right]^{\gamma}$ (9)
where $G$ is sample size, $N^{(a)}=\sum_{n^{\prime}}A_{n^{\prime}n^{\prime}}$ when $A_{nn}=a$, and $(K+G)N^{(a)}$ is the number of functional parameters in $P$ and $E$. During the tempering phase of the sampler, tempering is also applied to the sparsity-inducing term to avoid early convergence on a low rank before signatures are learned. Again, no tempered samples are used for posterior inference.
The structure of this prior results in a BIC-like penalty on rank. Specifically, once tempering is complete, the Gibbs update in Equation 3.2 is identical to the product of the standard BFI prior and the term $e^{-\frac{1}{2}\text{BIC}}$:
$\text{BIC}=-2\cdot\text{log}\left[p_{\text{Poisson}}(M|A_{nn}=a,...)\right]+\text{log}(G)\cdot(K+G)N^{(a)},$
$q^{a}(1-q)^{1-a}\cdot e^{-\frac{1}{2}\text{BIC}}=q^{a}(1-q)^{1-a}\cdot p(M|A_{nn}=a,...)\cdot G^{-\frac{1}{2}(K+G)N^{(a)}}.$